# Chow functions for partially ordered sets

### v1

@@ 1.1. Overview
Following another influential paper by Brenti [Bre99], we will refer to this as the Kazhdan–Lusztig–Stanley (KLS) theory for posets. We point to a recent survey by Proudfoot [Pro18] for a self-contained introduction to KLS theory and its algebro-geometric consequences. In what follows we summarize the basic setup of the KLS theory, following closely the notation of [Pro18]. In Section 2 we provide more detail about the construction of the main objects.
@@
The central contribution of the present work is the introduction of a new class of functions, that we call Chow functions, associated to any $(P,\rho)$-kernel $\kappa$. As opposed to the case of the KLS functions where a convention of left versus right constitutes an essential part of the definition, in our case there is a single distinguished element $\mathrm{H}\in\mathcal{I}_{\rho}(P)$ called the $\kappa$-Chow function associated to $(P,\rho)$. Notably, the KLS functions are required to satisfy a very restrictive degree bound: $\deg f_{st}<\frac{1}{2}\rho_{st}$ and $\deg g_{st}<\frac{1}{2}\rho_{st}$ for each $s<t$. In our case, the Chow function $\mathrm{H}$ satisfies a weaker degree bound: $\deg\mathrm{H}_{st}<\rho_{st}$ for each $s<t$ but, in order to compensate the additional degrees of freedom, one imposes that the polynomials $\mathrm{H}_{st}$ are palindromic.
As we will demonstrate in this paper, Chow functions and KLS functions are tightly related to each other. Often, properties of one have an impact on the other. The most significant example of this phenomenon in the present paper is the following result.
@@
  • Our framework does not require us to work with matroids nor posets with characteristic polynomials displaying any specific sign pattern in their coefficients, see Section 4. More so, we can apply these constructions to the examples of face lattices of polytopes (Section 5) or Bruhat intervals (Section 6).
@@
  • This framework is amenable to build upon intuition from one setting (say, polytopes) and use it in another one (say, Coxeter groups). For example, the use of the $\mathbf{cd}$-index in the case of polytopes in Section 5 led us to consider in Section 6 the complete $\mathbf{cd}$-index of Bruhat intervals introduced by Billera–Brenti [BB11].
@@ 1.2. The three main examples
A priori, we do not require the poset $P$ to be bounded. However, in some important cases we are in this situation and will correspondingly denote $\widehat{0}=\min P$ and $\widehat{1}=\max P$. We will refer to $\mathrm{H}_{\widehat{0}\widehat{1}}(x)$ as the Chow polynomial of $P$.
An important feature of Chow functions $\mathrm{H}$ is that they are symmetric. More precisely, each of the polynomials $\mathrm{H}_{st}(x)$ satisfies the identity
@@ 1.2.1. Characteristic Chow functions
As is pointed out in [Pro18], the characteristic function $\chi\in\mathcal{I}_{\rho}(P)$ is a $P$-kernel in any weakly ranked locally finite poset $P$, and lattices of flats of matroids are just a special case. In particular, there is no formal obstruction to consider the KLS functions $f$ and $g$ and the Chow function $\mathrm{H}$ arising from this setup. For the sake of clarity, we will refer to this Chow function $\mathrm{H}$ as the characteristic Chow function or, for brevity, the $\chi$-Chow function of $(P,\rho)$. In the matroid setting, one has the following result.
@@
The first part of the above statement is proved in our prequel [FMSV24], whereas the second follows from the validity of the Hard Lefschetz theorem, proved by Adiprasito, Huh, and Katz [AHK18].
In [FMSV24] we proved a strengthening of unimodality in the above statement: the Hilbert–Poincaré series of a matroid Chow ring is in fact $\gamma$-positive [FMSV24, Theorem 1.8]. The main tool to prove that was a key result of Braden, Huh, Matherne, Proudfoot, and Wang [BHM+22a], who established a semi-small decomposition for the Chow ring of a matroid.
@@
Besides unimodality, one may consider the stronger property of being $\gamma$-positive. For geometric lattices this property is known to hold true thanks to [FMSV24, Theorem 1.8]. We go far beyond geometric lattices and prove the following.
@@
For a general Cohen–Macaulay poset there is no obvious way of defining the Chow ring and therefore no clear analogue of the semi-small decomposition of [BHM+22a]. The last statement generalizes a beautiful result by Stump [Stu24, Theorem 1.1], which was a key motivation for our proof. In [FS24, Conjecture 8.18], Ferroni and Schröter conjectured that whenever $P=\mathcal{L}(\mathsf{M})$ is the lattice of flats of a matroid $\mathsf{M}$, then the Hilbert–Poincaré series of the Chow ring of $\mathsf{M}$ is a real-rooted polynomial. We formulate the stronger conjecture that this property also holds true for the $\chi$-Chow polynomials of all Cohen–Macaulay posets. Proving our conjecture would also imply another conjecture by Huh on the real-rootedness of Hilbert series of augmented Chow rings of matroids.
@@ 1.2.2. Eulerian Chow functions
Whenever $P$ is an Eulerian poset, the element $\varepsilon\in\mathcal{I}_{\rho}(P)$ given by $\varepsilon_{st}=(x-1)^{\rho_{st}}$ is a $P$-kernel. The resulting Chow function will be customarily called the Eulerian Chow function, or $\varepsilon$-Chow function for brevity, associated to $P$. We prove the following result.
@@
By barycentric subdivision of a poset $P$ we mean the simplicial complex whose faces are the flags of elements of $P$. We do not know whether Eulerian Chow polynomials are always non-negative. Moreover, we explain why we expect this question to be very subtle. By the positivity of the KLS functions proved in certain special cases (e.g., for face posets of simplicial polytopes [Sta80], of general polytopes [Kar04], or of simplicial spheres [Adi18, PP20]), Theorem 3.12 guarantees that the $\varepsilon$-Chow function is non-negative and unimodal. However, another deep result by Karu [Kar06] about the $\mathbf{cd}$-index of Gorenstein* posets (that is, posets that are both Eulerian and Cohen–Macaulay) can be used to obtain the following stronger property.
@@ 1.2.3. Coxeter Chow functions
In the above statement $B(u,v)$ stands for all the paths in the Bruhat graph of $W$ that go from $u$ to $v$, $\ell(\Delta)$ stands for the length of the path $\Delta$, $\operatorname{des}$ stands for the number of descents of the path, whereas asc stands for the number of ascents. In particular, the Chow function is enumerating these paths according to a descent-like statistic. We show that the combinatorial invariance conjecture for Chow functions is equivalent to the combinatorial invariance conjecture for Kazhdan–Lusztig or $R$-polynomials, see Theorem 6.16.
@@ 1.3. Paper outline
The chief contribution of this paper is the combinatorial framework of Chow functions developed in Section 3; we view this construction as a counterpart for Stanley’s development of the theory of KLS functions in [Sta92]. This section introduces Chow functions and augmented Chow functions, and here we study their general connection with KLS functions and $Z$-functions. In this section we prove various numerical analogues of graded module decompositions appearing in [BHM+22b].
Section 4 comprises the first central example of how to apply the machinery developed in Section 3: we introduce the $\chi$-Chow function and study a number of combinatorial properties that it and its augmented counterparts satisfy; furthermore, we use the case of matroids to explain the algebro-geometric motivation for the main results in Section 3.
@@ 2.1. Inequalities for polynomials
Let $p(x)=a_{0}+a_{1}x+\cdots+a_{m}x^{m}$ denote a polynomial having non-negative coefficients. The polynomial $p(x)$ is said to be unimodal if there exists an index $j$ such that
@@
The proof is straightforward so we omit it. A further property that will be of relevance in the present paper is that of $\gamma$-positivity. We say that the polynomial $p(x)$ is $\gamma$-positive if it is symmetric with center of symmetry $d$ and there exist non-negative integers $\gamma_{0},\ldots,\gamma_{\lfloor d/2\rfloor}$ such that
@@ 2.2. Essential notions about posets
    There is a multiplicative identity in $\mathcal{I}(P)$, denoted $\delta\in\mathcal{I}(P)$ and defined by
@@ 2.3. Weak rank functions and characteristic functions
A weak rank function $\rho$ on a locally finite poset $P$ induces a special subalgebra $\mathcal{I}_{\rho}(P)\subseteq\mathcal{I}(P)$, defined by
@@
This subalgebra admits an involution $a\mapsto a^{\operatorname{rev}}$ defined via the following identity:
@@ 2.4. The basics of KLS theory
Let $P$ be the Boolean lattice having $3$ atoms, and regard it as a graded poset so that $\rho_{st}$ is the length of any saturated chain starting at $s$ and ending at $t$. Fix any number $m\in\mathbb{Z}$ and define the following element $\kappa$ on $\mathcal{I}_{\rho}(P)$:
@@ 2.5. Graded posets, flag $f$-vectors, and Cohen–Macaulayness
$\alpha_{P}(S)=\{\text{chains $w_{1}<\cdots<w_{m}$ in $P$}:\rho(w_{i})=s_{i}\text{ for $1\leq i\leq m$}\}.$
The map $\alpha\colon 2^{\{0,\ldots,r\}}\to\mathbb{Z}_{\geq 0}$ is commonly known as the flag $f$-vector of $P$. It can be encoded in an alternative way by considering the flag $h$-vector, which is the map $\beta\colon 2^{\{0,\ldots,r\}}\to\mathbb{Z}$ defined by the condition:
@@
Let $\Delta$ be a simplicial complex, and denote by $f_{i}$ the number of faces of $\Delta$ having cardinality [^3] We are using the conventions of Björner in [Bjö92]. $i$ (or, equivalently, dimension $i-1$). The $f$-vector of $\Delta$ is defined by
$f(\Delta)=(f_{0},\ldots,f_{d-1})$
where $d=\dim(\Delta)$. The $f$-polynomial of $\Delta$ is the polynomial $f(\Delta,x)=f_{0}x^{d}+f_{1}x^{d-1}+\cdots+f_{d}$. The $h$-vector and the $h$-polynomial of $\Delta$ are defined via
@@
Recall that to every poset $P$ we associate a simplicial complex $\Delta(P)$, called the order complex of $P$. The faces of $\Delta(P)$ correspond to chains of elements in $P$. The $f$-vector of the simplical complex $\Delta(P)$ is encoded in the flag $f$-vector of $P$. Put precisely,
@@
It is not difficult to show that $h$-vector of $\Delta(P)$, is given by
@@
Note that the flag $f$-vector of $P$ has non-negative values as it count chains. However, the flag $h$-vector often fails to be non-negative. Similarly, the $h$-vector of $\Delta(P)$ can a priori have negative coefficients. In what follows we recapitulate an important case in which the flag $h$-vector of $P$ is indeed non-negative (and therefore the $h$-vector of $\Delta(P)$).
As any other simplicial complex, $\Delta(P)$ admits a geometric realization, that we will denote $|\Delta(P)|$. Note that every (open) interval $(s,t)=\{w\in P:s<w<t\}$ is itself a graded poset. By definition, we say that $P$ is Cohen–Macaulay (over $\mathbb{Q}$) if the rational reduced homology groups of the order complex of every open interval $(s,t)$ satisfy
@@
In other words, the (reduced) homology of every open interval $(s,t)$ must be concentrated in dimension $\dim\Delta(s,t)=\rho_{st}-2$. The class of Cohen–Macaulay posets comprises a number of well-studied families, such as distributive lattices, posets that admit an R-labelling, posets that are EL-shellable, etc. It is worth noting that Cohen–Macaulayness is a topological property, that is, if $P_{1}$ and $P_{2}$ are posets and there is an homeomorphism $|\Delta(P_{1})|\approx|\Delta(P_{2})|$ then $P_{1}$ is Cohen–Macaulay if and only if $P_{2}$ is Cohen–Macaulay.
@@
Let $P$ a finite, graded, bounded poset. If $P$ is Cohen–Macaulay, then the flag $h$-vector of $P$ has non-negative entries.
@@ 3.1. Reduced kernels
A well-known object in the theory of hyperplane arrangements is the “reduced” characteristic polynomial. If $\mathcal{A}$ denotes a non-empty central hyperplane arrangement, the characteristic polynomial $\chi_{\mathcal{A}}(x)$ vanishes when evaluated at $x=1$ (this is an immediate fact that follows from the definition of the Möbius function). In this setting, and in the more general context of matroids, it is customary to the define the reduced characteristic polynomial by $\overline{\chi}_{\mathcal{A}}(x)=\frac{1}{x-1}\,\chi_{\mathcal{A}}(x)$. We point out that this agrees with the usual notation and conventions used, for instance, in [AHK18, Definition 9.1].
@@
$\kappa_{st}(x)+\kappa_{st}^{\operatorname{rev}}(x)=-\sum_{s<w<t}\kappa_{sw}(x)\kappa_{wt}^{\operatorname{rev}}(x),$
which implies again that $\kappa_{st}(x)+\kappa_{st}^{\operatorname{rev}}(x)=0$, as all the polynomials on the right hand side are associated to smaller intervals that are not singletons. ∎
@@
The choice we impose on defining $\kappa_{ss}(x)$ as $-1$ (as opposed to just $1$) is in fact important. An alternative approach would be to define $\overline{\kappa}_{st}(x)$ as $\frac{1}{1-x}\kappa_{st}(x)$ for $s<t$ and as $1$ for $s=t$, but this would create confusions with the standard notation for the reduced characteristic polynomial. Since the latter is a key motivating example, we prefer to follow the convention on the statement of Definition 3.2. On the other hand, we warn the reader that in other sources the notation $\overline{\kappa}$ stands for what we denote here as $\kappa^{\operatorname{rev}}$. We prefer to use the overline to denote “reduced” instead of “reversed”.

### v2

@@ 1.1. Overview
Following another influential paper by Brenti [Bre99], we will refer to this as the Kazhdan–Lusztig–Stanley (KLS) theory for posets. We point to a recent survey by Proudfoot [Pro18] for a self-contained introduction to KLS theory and its algebro-geometric consequences. In what follows we summarize the basic setup of the KLS theory, following closely the notation of [Pro18]. In Section 2 we provide more details about the construction of the main objects.
@@
The central contribution of the present work is the introduction of a new class of functions, that we call Chow functions, associated to any $(P,\rho)$-kernel $\kappa$. As opposed to the case of the KLS functions, where a convention of left versus right constitutes an essential part of the definition, in our case there is a single distinguished element $\mathrm{H}\in\mathcal{I}_{\rho}(P)$ called the $\kappa$-Chow function associated to $(P,\rho)$. Notably, the KLS functions are required to satisfy a very restrictive degree bound: $\deg f_{st}<\frac{1}{2}\rho_{st}$ and $\deg g_{st}<\frac{1}{2}\rho_{st}$ for each $s<t$. In our case, the Chow function $\mathrm{H}$ satisfies a weaker degree bound: $\deg\mathrm{H}_{st}<\rho_{st}$ for each $s<t$; in order to compensate for the additional degrees of freedom, one imposes that the polynomials $\mathrm{H}_{st}$ are palindromic.
Chow functions and KLS functions turn out to be tightly connected to each other. Often, properties of one have an impact on the other. The most significant example of this phenomenon in the present paper is the following result.
@@
  • Our framework does not require us to work with matroids nor posets with characteristic polynomials displaying any specific sign pattern in their coefficients, see Section 4. In particular, we can apply these constructions to the examples of face lattices of polytopes (Section 5) or Bruhat intervals (Section 6).
@@
  • This framework is amenable for building upon intuition from one setting (say, polytopes) and using it in another (say, Coxeter groups). For example, the use of the $\mathbf{cd}$-index in the case of polytopes in Section 5 led us to consider in Section 6 the complete $\mathbf{cd}$-index of Bruhat intervals introduced by Billera–Brenti [BB11].
@@ 1.2. The three main examples
A priori, we do not require the poset $P$ to be bounded. However, in some important cases we find ourselves in this situation and will correspondingly denote $\widehat{0}=\min P$ and $\widehat{1}=\max P$. We will refer to $\mathrm{H}_{\widehat{0}\widehat{1}}(x)$ as the Chow polynomial of $P$.
An important feature of Chow functions is that they are symmetric. More precisely, each of the polynomials $\mathrm{H}_{st}(x)$ satisfies the identity
@@ 1.2.1. Characteristic Chow functions
As is pointed out in [Pro18], the characteristic function $\chi\in\mathcal{I}_{\rho}(P)$ is a $P$-kernel in any weakly ranked locally finite poset $P$, and lattices of flats of matroids are just a special case. In particular, there is no formal obstruction in considering the KLS functions $f$ and $g$ and the Chow function $\mathrm{H}$ arising from this setup. For the sake of clarity, we will refer to this Chow function $\mathrm{H}$ as the characteristic Chow function or, for brevity, the $\chi$-Chow function of $(P,\rho)$. In the matroid setting, one has the following result.
@@
The first part of the above statement is proved in [FMSV24], whereas the second follows from the validity of the Hard Lefschetz theorem, proved by Adiprasito, Huh, and Katz [AHK18].
In [FMSV24] a strengthening of unimodality is also proved: the Hilbert–Poincaré series of a matroid Chow ring is in fact $\gamma$-positive [FMSV24, Theorem 1.8]. The main tool to prove that was a key result of Braden, Huh, Matherne, Proudfoot, and Wang [BHM+22a], who established a semi-small decomposition for the Chow ring of a matroid.
@@
Besides unimodality, one may consider the stronger property of being $\gamma$-positive. For geometric lattices this property is known to hold via [FMSV24, Theorem 1.8], but in general it does not (see Example 4.19 below). We go far beyond geometric lattices and prove the following.
@@
For a general Cohen–Macaulay poset there is no obvious way of defining the Chow ring and therefore no clear analogue of the semi-small decomposition of [BHM+22a]. In fact, we prove a statement more general than the above result. We show that the $\gamma$-polynomial associated to the $\chi$-Chow function is a non-negative specialization of the flag $h$-vector (see Theorem 4.25). This also generalizes a beautiful result by Stump [Stu24, Theorem 1.1], which was a key motivation for our proof. In [FS24, Conjecture 8.18], Ferroni and Schröter conjectured that whenever $P=\mathcal{L}(\mathsf{M})$ is the lattice of flats of a matroid $\mathsf{M}$, then the Hilbert–Poincaré series of the Chow ring of $\mathsf{M}$ is a real-rooted polynomial. We formulate the stronger conjecture that this property also holds true for the $\chi$-Chow polynomials of all Cohen–Macaulay posets. Proving our conjecture would also imply another conjecture by Huh on the real-rootedness of Hilbert series of augmented Chow rings of matroids.
@@ 1.2.2. Eulerian Chow functions
Whenever $P$ is an Eulerian poset, the element $\varepsilon\in\mathcal{I}_{\rho}(P)$ given by $\varepsilon_{st}=(x-1)^{\rho_{st}}$ is a $P$-kernel. The resulting Chow function will be called the Eulerian Chow function, or $\varepsilon$-Chow function for brevity, associated to $P$. We prove the following result.
@@
By barycentric subdivision of a poset $P$, we mean the simplicial complex whose faces are the flags of elements of $P$. We do not know whether Eulerian Chow polynomials are always non-negative. Moreover, we explain why we expect this question to be very subtle. By the positivity of the KLS functions proved in certain special cases (e.g., for face posets of simplicial polytopes [Sta80], of general polytopes [Kar04], or of simplicial spheres [Adi18, PP20]), Theorem 1.1 guarantees that the $\varepsilon$-Chow function is non-negative and unimodal. However, another deep result by Karu [Kar06] about the $\mathbf{cd}$-index of Gorenstein* posets (that is, posets that are both Eulerian and Cohen–Macaulay) can be used to obtain the following stronger property.
@@ 1.2.3. Coxeter Chow functions
Let $W$ be a Coxeter group with a reflection order $<$ and two elements $u,v\in W$. Then
@@
In the above statement $B(u,v)$ stands for all the paths in the Bruhat graph of $W$ that go from $u$ to $v$, $\ell(\Delta)$ stands for the length of the path $\Delta$, $\operatorname{des}$ stands for the number of descents of the path, whereas asc stands for the number of ascents. In particular, the Chow function enumerates these paths according to a descent-like statistic. We show that the combinatorial invariance conjecture for Chow functions is equivalent to the combinatorial invariance conjecture for Kazhdan–Lusztig or $R$-polynomials, see Theorem 6.16.
@@ 1.3. Paper outline
The chief contribution of this paper is the combinatorial framework of Chow functions developed in Section 3; we view this construction as a counterpart to Stanley’s development of the theory of KLS functions in [Sta92]. This section introduces Chow functions and augmented Chow functions, and here we study their general connection with KLS functions and $Z$-functions. In this section we prove various numerical analogues of graded module decompositions appearing in [BHM+22b]. Sections 45, and 6 then apply the general machinery developed in Section 3 to our three central running examples, as we now describe.
In Section 4, we introduce the $\chi$-Chow function and study a number of combinatorial properties that it and its augmented counterparts satisfy; furthermore, we use the case of matroids to explain the algebro-geometric motivation for the main results in Section 3.
@@ 2.1. Inequalities for polynomials
Let $p(x)=a_{0}+a_{1}x+\cdots+a_{m}x^{m}\in\mathbb{Z}[x]$ denote a polynomial having non-negative coefficients. The polynomial $p(x)$ is said to be unimodal if there exists an index $j$ such that
@@
The proof is straightforward, so we omit it. A further property that will be of relevance in the present paper is that of $\gamma$-positivity. We say that the polynomial $p(x)$ is $\gamma$-positive if it is symmetric with center of symmetry $d/2$ and there exist non-negative integers $\gamma_{0},\ldots,\gamma_{\lfloor d/2\rfloor}$ such that
@@ 2.2. Essential notions about posets
    There is a multiplicative identity $\delta\in\mathcal{I}(P)$ defined by
@@ 2.3. Weak rank functions and characteristic functions
A weak rank function $\rho$ on a locally finite poset $P$ induces a special subalgebra $\mathcal{I}_{\rho}(P)\subseteq\mathcal{I}(P)$ defined by
@@
This subalgebra admits an involution $a\mapsto a^{\operatorname{rev}}$ defined via the identity
@@ 2.4. The basics of KLS theory
Let $P$ be the Boolean lattice having three atoms, and regard it as a graded poset so that $\rho_{st}$ is the length of any saturated chain starting at $s$ and ending at $t$. Fix any number $m\in\mathbb{Z}$ and define the following element $\kappa$ on $\mathcal{I}_{\rho}(P)$:
@@ 2.5. Graded posets, flag $f$-vectors, and Cohen–Macaulayness
$\alpha_{P}(S)=\|\{\text{chains $w_{1}<\cdots<w_{m}$ in $P$}:\rho(w_{i})=s_{i}\text{ for $1\leq i\leq m$}\}\|.$
The map $\alpha_{P}\colon 2^{\{0,\ldots,r\}}\to\mathbb{Z}_{\geq 0}$ is commonly known as the flag $f$-vector of $P$. It can be encoded in an alternative way by considering instead the flag $h$-vector, which is the map $\beta_{P}\colon 2^{\{0,\ldots,r\}}\to\mathbb{Z}$ defined by the condition
@@
Let $\Delta$ be a simplicial complex, and denote by $f_{i}$ the number of faces of $\Delta$ having cardinality $i$ (or, equivalently, dimension $i-1$) [^3] We are using the conventions of Björner in [Bjö92].. The $f$-vector of $\Delta$ is defined by
$f(\Delta)=(f_{0},\ldots,f_{d})$
where $d=\dim(\Delta)+1$. The $f$-polynomial of $\Delta$ is the polynomial $f(\Delta,x)=f_{0}x^{d}+f_{1}x^{d-1}+\cdots+f_{d}$. The $h$-vector and the $h$-polynomial of $\Delta$ are defined via
@@
Recall that to every poset $P$ we may associate a simplicial complex $\Delta(P)$, called the order complex of $P$. The faces of $\Delta(P)$ correspond to chains of elements in $P$. The $f$-vector of the simplicial complex $\Delta(P)$ is encoded in the flag $f$-vector of $P$. Put precisely,
@@
It is not difficult to show that the $h$-vector of $\Delta(P)$ is given by
@@
Note that the flag $f$-vector of $P$ has non-negative values, as it count chains. However, the flag $h$-vector often fails to be non-negative. Similarly, the $h$-vector of $\Delta(P)$ can a priori have negative coefficients. In what follows we recapitulate an important case in which the flag $h$-vector of $P$ is indeed non-negative (and therefore the $h$-vector of $\Delta(P)$).
As any other simplicial complex, $\Delta(P)$ admits a geometric realization that we will denote $|\Delta(P)|$. Note that every (open) interval $(s,t)=\{w\in P:s<w<t\}$ is itself a graded poset. By definition, we say that $P$ is Cohen–Macaulay (over $\mathbb{Q}$) if the rational reduced homology groups of the order complex of every open interval $(s,t)$ satisfy
@@
In other words, the (reduced) homology of every open interval $(s,t)$ must be concentrated in dimension $\dim\Delta(s,t)=\rho_{st}-2$. The class of Cohen–Macaulay posets comprises a number of well-studied families, such as distributive lattices, posets that admit an R-labelling, posets that are EL-shellable, etc. It is worth noting that Cohen–Macaulayness is a topological property, that is, if $P_{1}$ and $P_{2}$ are posets and there is a homeomorphism $|\Delta(P_{1})|\approx|\Delta(P_{2})|$ then $P_{1}$ is Cohen–Macaulay if and only if $P_{2}$ is Cohen–Macaulay.
@@
Let $P$ be a finite graded bounded poset. If $P$ is Cohen–Macaulay, then the flag $h$-vector of $P$ has non-negative entries.
@@ 3.1. Reduced kernels
A well-known object in the theory of hyperplane arrangements is the “reduced” characteristic polynomial. If $\mathcal{A}$ denotes a non-empty central hyperplane arrangement, the characteristic polynomial $\chi_{\mathcal{A}}(x)$ vanishes when evaluated at $x=1$. (This is an immediate fact that follows from the definition of the Möbius function.) In this setting, and in the more general context of matroids, it is customary to define the reduced characteristic polynomial by $\overline{\chi}_{\mathcal{A}}(x)=\frac{1}{x-1}\,\chi_{\mathcal{A}}(x)$. We point out that this agrees with the usual notation and conventions used, for instance, in [AHK18, Definition 9.1].
@@
$\kappa_{st}(x)+\kappa_{st}^{\operatorname{rev}}(x)=-\sum_{s<w<t}\kappa_{sw}(x)\kappa_{wt}^{\operatorname{rev}}(x).$
As all the polynomials on the right hand side are associated to smaller intervals that are not singletons, evaluating at $x=1$ gives $\kappa_{st}(1)+\kappa_{st}^{\operatorname{rev}}(1)=0$, which finishes the proof. ∎
@@
The choice we impose on defining $\kappa_{ss}(x)$ as $-1$ (as opposed to $1$) is intentional. An alternative approach would be to define $\overline{\kappa}_{st}(x)$ as $\frac{1}{1-x}\kappa_{st}(x)$ for $s<t$ and as $1$ for $s=t$, but this might cause confusion with the standard notation for the reduced characteristic polynomial. Since the latter is a key motivating example, we prefer to follow the convention in Definition 3.2. On another note, we warn the reader that in other sources the notation $\overline{\kappa}$ stands for what we denote here as $\kappa^{\operatorname{rev}}$. We prefer to use the overline to denote “reduced” instead of “reversed”.